HETAV PATEL

Grade 11,

The specific heats of the gas are Cp=0.2cal/gm°C and Cv=0.15cal/gm°C find a) the molar mass of the gas b) the number of degree of freedom of the gas molecules [ Take R=2cal/mol.°C]

The specific heats of the gas are Cp=0.2cal/gm°C and Cv=0.15cal/gm°C find 
a) the molar mass of the gas
b) the number of degree of freedom of the gas molecules                       [ Take R=2cal/mol.°C]
 

Grade:11

1 Answers

Arun
5 years ago
Dear Hemant
 
As we know.
Cp – Cv = n* R
0.2 – 0.15 = n * 0.0821
0.05/0.0821 = n
n = 0.61 moles
 
Regards
Arun